OpenStudy (anonymous):

Simplify the expression. x times the square root of the quantity 64 x cubed, plus 2 x times the square root of x, plus 8 times the square root of the quantity 2 x If the simplified expression is written in standard form, what is the leading coefficient?

OpenStudy (anonymous):

hold on I'm doing it lol

OpenStudy (agent0smith):

\[\Large x \sqrt{64x^3} + 2x \sqrt{x} + 8 \sqrt{2x}\]?

OpenStudy (anonymous):

\[x \sqrt{64x^3} + 2x \sqrt{x} + 8\sqrt{2x}\]

OpenStudy (agent0smith):

This will help you get started\[\Large x \sqrt{64} \sqrt{x^3} + 2x \sqrt{x} + 8 \sqrt{2x}\]x^3 is x^2x:\[\Large x \sqrt{64} \sqrt {x^2 \times x} + 2x \sqrt{x} + 8 \sqrt{2x}\] \[\Large x \sqrt{64} \sqrt {x^2 }\sqrt{x} + 2x \sqrt{x} + 8 \sqrt{2x}\]

OpenStudy (anonymous):

x 8 sqrt x^2 sqrt x + 2x sqrt x + 8 sqrt 2x

OpenStudy (agent0smith):

So now you have\[\Large 8x \sqrt {x^2 }\sqrt{x} + 2x \sqrt{x} + 8 \sqrt{2x}\] and what's the sqrt of x^2?

OpenStudy (anonymous):

2x

OpenStudy (agent0smith):

Remember we went over this a few days ago. Sqrt of 4^2 is 4. Sqrt of 2^2 is 2. Sqrt of x^2 is...

OpenStudy (anonymous):

oh x

OpenStudy (agent0smith):

\[\Large 8x \times x\sqrt{x} + 2x \sqrt{x} + 8 \sqrt{2x}\]

OpenStudy (anonymous):

now what do I do?

OpenStudy (agent0smith):

Simplify that, and combine like terms

OpenStudy (agent0smith):

x times x is...?

OpenStudy (anonymous):

2x

OpenStudy (agent0smith):

x + x is 2x. x*x is not 2x. 4*4 = 16, which is the same as 4^2. x*x = ?

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

I think I got it tell me if im right, 16x x 2x+3x 16x x 5x 21x?

OpenStudy (anonymous):

that's not right. it has to be one of the choices above :(

OpenStudy (anonymous):

are you gonna help me @agent0smith ?

OpenStudy (agent0smith):

Sorry, i gotta go.

OpenStudy (anonymous):

thanks for the help :(

OpenStudy (agent0smith):

\[\Large 8x ^2\sqrt{x} + 2x \sqrt{x} + 8 \sqrt{2x}\]

OpenStudy (agent0smith):

@jaderbrown still need to finish this...?

OpenStudy (agent0smith):

You can simplify a bit more depending on the answer choices

OpenStudy (anonymous):

I tried to do it a few hours ago but I don't know if I got the right answer

OpenStudy (anonymous):

and I don't remember what I put in for the answer